# When Attention Sink Emerges in Language Models: An Empirical View

### v1

@@ B.2 Attention sink under different data domains
B.3 More visualizations of internal mechanism of attention sink in LMs
$\ell_{2}$-norm. We first show that large $\ell_{2}$-norm of hidden states $\bm{h}_{1}^{l}$ and small $\ell_{2}$-norm of keys $\bm{k}_{1}^{l\textrm{,}\,h}$ and values $\bm{v}_{1}^{l\textrm{,}\,h}$ (especially for values) universally exist in open-sourced LMs, including LLaMA2-7B Base (Figure 10), GPT2-Large (Figure 11), Mistral-7B (Figure 12), and Pythia-1B (Figure 13). We also showcase our small-sized LLaMA model under the default setup in Figure 14. It is noted that for the final transformer block $l=L$, we take the hidden states before LN. We note that different LMs may have different starting blocks where massive activations appear.
@@ B.3 More visualizations of internal mechanism of attention sink in LMs
Figure 14: $\ell_{2}$-norm of hidden states/keys/values of the first token/other tokens in our small LLaMA model.
@@
Figure 20: Distribution of importance scores for the first token across different blocks and heads in the LLaMA3/LLaMA3.1 family.
Figure 21: Distribution of importance scores for the first token across different blocks and heads in the Mistral family.
Figure 22: Distribution of importance scores for the first token across different blocks and heads in the GPT2 family.
Figure 23: Distribution of importance scores for the first token across different blocks and heads in the Pythia family.
Figure 24: Distribution of importance scores for the first token across different blocks and heads in the OPT family.
B.4 Huggingface links for open-sourced LMs
@@
Appendix C More experiments in LM pre-training
C.1 Optimization
@@
C.2 Data distribution
@@ C.4 Attention design
Therefore, this normalization scale can be regarded as the scale for $\bm{W}_{V}$ or $\bm{W}_{O}$. We show that this normalization scaling could be implemented by scaling learning rates and initialization. We use the $s$ to represent the optimization step, and $s=\textrm{0}$ refers to the initialization. When scaling the normalization, we have the following update rule (take $\bm{W}_{O}$ for example):
$\bm{W}_{O}^{s+1}$ $=\bm{W}_{O}^{s}-\eta\nabla\mathcal{L}(\alpha\bm{W}_{O}^{s})$ (32)
$=\bm{W}_{O}^{s}-\alpha\eta\nabla\mathcal{L}(\bm{W})|_{\bm{W}=\alpha\bm{W}_{O}^{s}}\textrm{,}$ (33)
where $\eta$ is the original learning rate. Suppose we only modify the learning rate and initialization, to ensure each optimization step $\hat{\bm{W}}_{O}^{s}=\bm{W}_{O}^{s}$, we need to first ensure $\hat{\bm{W}}_{O}^{0}=\alpha\bm{W}_{O}^{0}$. Suppose that we have $\hat{\bm{W}}_{O}^{s}=\bm{W}_{O}^{s}$, then the update rule is:
$\hat{\bm{W}}_{O}^{s+1}$ $=\hat{\bm{W}}_{O}^{s}-\eta^{\prime}\nabla\mathcal{L}(\hat{\bm{W}}_{O}^{s})$ (34)
$=\alpha\bm{W}_{O}^{s}-\eta^{\prime}\nabla\mathcal{L}(\bm{W})|_{\bm{W}=\alpha\bm{W}_{O}^{s}}\textrm{,}$ (35)
To ensure $\hat{\bm{W}}_{O}^{s+1}=\bm{W}_{O}^{s+1}$, we need the new learning rate $\eta^{\prime}$ meets $\eta^{\prime}=\alpha^{2}\eta$.
@@
$\textrm{exp}(\frac{\bm{q}_{i}^{T}\bm{k}_{j}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{T}}{\sqrt{d_{h}}})$ 18.18 3.73
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}})$ $1$       0.44 3.70
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{T}}{\sqrt{d_{h}}})$ 30.24 3.74
$\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}})+1$ $1$       0.80 3.69
$\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}})+1$ $\sum_{j^{\prime}=1}^{i}\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{T}}{\sqrt{d_{h}}})+1$
$\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j})+1)^{T}}{\sqrt{d_{h}}}$ $\sum_{j^{\prime}=1}^{i}\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j^{\prime}})+1)^{T}}{\sqrt{d_{h}}}$    53.65 4.19
$\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j})+1)^{T}}{\sqrt{d_{h}}}$ $1$
$\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}}$ $\textrm{max}\left(\left|\sum_{j^{\prime}=1}^{i}\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{T}}{\sqrt{d_{h}}}\right|\textrm{,}\,1\right)$
$\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}}$ $1$    0.00 3.99
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{T}}{\sqrt{d_{h}}}$ $\textrm{max}\left(\left|\sum_{j^{\prime}=1}^{i}\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j^{\prime}})^{T}}{\sqrt{d_{h}}}\right|\textrm{,}\,1\right)$    0.19 3.85
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{T}}{\sqrt{d_{h}}}$ $1$    0.74 3.91

### v2

@@ C.2 Attention sink under different data domains
C.3 Attention sink under different pre-trained LMs
Relation to LM performance. We leverage the platform (Gao et al., 2024) to evaluate the performance of open-sourced LMs, including LLaMA2/LLaMA3/OPT/Pythia/GPT2 families, on downstream LM task, e.g., HellaSwag (Zellers et al., 2019). The results are visualized parallel with our attention sink metric in Figure 10, including both accuracy (Acc) and accuracy under normalization (Acc_Norm). We find that within the same LM family, with the increase of model scale, both attention sink amplitude and downstream LM performance are increasing. However, across different LM families, stronger attention sink does not always correlate with better performance. For instance, the OPT family has stronger attention sink than Pythia under the comparable model scale. However, the downstream LM performance is comparable.
Figure 10: Attention sink and downstream performance for various pre-trained LMs.
$\ell_{2}$-norm. We first show that large $\ell_{2}$-norm of hidden states $\bm{h}_{1}^{l}$ and small $\ell_{2}$-norm of keys $\bm{k}_{1}^{l\textrm{,}\,h}$ and values $\bm{v}_{1}^{l\textrm{,}\,h}$ (especially for values) universally exist in open-sourced LMs, including LLaMA2-7B Base (Figure 11), GPT2-Large (Figure 12), Mistral-7B (Figure 13), and Pythia-1B (Figure 14). It is noted that for the final transformer block $l=L$, we take the hidden states before LN. We note that different LMs may have different starting blocks where massive activations appear.
@@ C.3 Attention sink under different pre-trained LMs
Figure 20: Distribution of importance scores for the first token across blocks and heads in the LLaMA3/LLaMA3.1 family.
Figure 21: Distribution of importance scores for the first token across blocks and heads in the Mistral family.
Figure 22: Distribution of importance scores for the first token across blocks and heads in the GPT2 family.
Figure 23: Distribution of importance scores for the first token across blocks and heads in the Pythia family.
Jamba. Besides the auto-regressive Transformers, we also consider Jamba (Lieber et al., 2024; Team et al., 2024), a new foundation language model. Both Jamba-v0.1 (Lieber et al., 2024) and Jamba-1.5 Mini (Team et al., 2024) have 4 Jamba blocks, each of which includes 3 Mamba layers (Gu & Dao, 2023), 4 Mamba MoE layers (Shazeer et al., 2017; Fedus et al., 2022), and 1 Transformer layer. This adds to 32 layers (including 4 Transformer attention layers), 52B available parameters, and 12B active parameters in total. Firstly, we evaluate the attention sink metric and find that $\textrm{Sink}_{1}^{\epsilon}=\textrm{88.48}\%$ for Jamba-v0.1 and $\textrm{Sink}_{1}^{\epsilon}=\textrm{87.88}\%$ for Jamba-1.5 Mini, which indicates a strong attention sink on the first token. Then we visualize attention scores in several heads, as shown in Figure 25 and Figure 26. We also visualize the distribution of importance scores for the first token across blocks and heads in Jamba models in Figure 27. We observe that most heads have obvious attention sink, except for several heads in the 3rd Transformer layer.
Figure 24: Distribution of importance scores for the first token across blocks and heads in the OPT family.
Figure 25: Attention sink in Jamba-v0.1.
Figure 26: Attention sink in Jamba-1.5 Mini.
Figure 27: Distribution of importance scores for the first token across blocks and heads in the Jamba models.
C.4 Huggingface links for open-sourced LMs
@@
Jamba-v0.1 ai21labs/Jamba-v0.1
Jamba-1.5 Mini ai21labs/AI21-Jamba-1.5-Mini
Appendix D More experiments in LM pre-training
D.1 Optimization
Learning rate. In Section 4, we find that attention sink appears less obvious in LMs trained with small learning rates. This conclusion holds even if we compensate for more training steps. In our basic setup, we adopt a learning rate of 4e-4 for 20k training steps. When we scale the learning rate to half, i.e., 2e-4, we scale the training steps to 2 times, i.e., 40k. As shown in Table 9, when we keep the multiply between learning rate and training steps constant (highlighted using cyan color), LMs trained with smaller learning rates tend to exhibit less obvious attention sink. Therefore, we conclude that small learning rates not only slow down the increase of attention sink but also mitigate attention sink even with longer training durations.
Table 9: Attention sink appears less obvious in LMs trained with small learning rates even compensating for more training steps.
learning rate training steps (k) $\textrm{Sink}_{1}^{\epsilon}$(%) valid loss
8e-4 10 23.44 3.79
8e-4 20 32.23 3.70
4e-4 20 18.18 3.73
2e-4 20 11.21 3.78
2e-4 40 16.81 3.68
1e-4 20    2.90 3.92
1e-4 80    6.29 3.67
@@ D.1 Optimization
D.2 Data distribution
Training data amount. In Section 5, we find that with less training data, attention sink also disappears. Meanwhile, LMs are also prone for overfitting. To disentangle the effects of training data amount and overfitting on attention sink, we monitor the dynamics of train/valid loss and attention sink metric during the LM pre-training in a more granular level, as present in Figure 28. With only 50M and 100M training data, LMs overfit at very early stages, between 1k and 2k steps. Meanwhile, $\textrm{Sink}_{1}^{\epsilon}$ maintains a very small value (less than 1%). While for the setup of 5B training data, $\textrm{Sink}_{1}^{\epsilon}$ keeps increasing after a certain step. This indicates that the amount of training data, instead of overfitting, plays an important role in the emergence of attention sink.
Figure 28: Dynamics of train/valid loss and $\textrm{Sink}_{1}^{\epsilon}$ during LM pre-training under different amounts of training data: (Left) 50M; (Middle) 100M; (Right) 5B.
@@ D.4 Attention design
Table 13: Even with very few learnable dimensions for $\bm{k}^{*l\textrm{,}h}$, large attention appears in $\bm{k}^{*l\textrm{,}h}$.
$d_{a}$ 1 2 4 8 16 32 64
$\textrm{Sink}_{1}^{\epsilon}$(%)    4.74    4.96    4.39    4.54    2.19    1.94    0.04
valid loss    3.73    3.72    3.72    3.73    3.73    3.73    3.72
Learnable dimensions of K biases. In the setup of K biases, we set all dimensions of $\bm{k}^{*l\textrm{,}\,h}$ as learnable weights. In Section 3.1, we have shown that K biases are distributed in a different manifold, with low rank. Therefore, we consider only $d_{a}$ dimensions of $\bm{k}^{*l\textrm{,}\,h}$ are adjustable/learnable while the other $d_{h}-d_{a}$ dimensions are zeros. As present in Table 13, with very few learnable dimensions, even for $d_{a}=\textrm{1}$, $\bm{k}^{*l\textrm{,}\,h}$ are still allocated significant attention. With more learnable dimensions, the attention sink appears more obvious.
@@
Therefore, this normalization scale can be regarded as the scale for $\bm{W}_{V}$ or $\bm{W}_{O}$. We show that this normalization scaling could be implemented by scaling learning rates and initialization. We use the $s$ to represent the optimization step, and $s=\textrm{0}$ refers to the initialization. When scaling the normalization, we have the following SGD update rule (take $\bm{W}_{O}$ for example):
$\bm{W}_{O}^{s+1}$ $=\bm{W}_{O}^{s}-\eta\nabla_{\bm{W}_{O}^{s}}\mathcal{L}(\alpha\bm{W}_{O}^{s})$ (32)
$=\bm{W}_{O}^{s}-\alpha\eta\nabla_{\bm{W}}\mathcal{L}(\bm{W})|_{\bm{W}=\alpha\bm{W}_{O}^{s}}\textrm{,}$ (33)
where $\eta$ is the original learning rate. Suppose we only modify the learning rate and initialization, to ensure each optimization step $\hat{\bm{W}}_{O}^{s}=\alpha\bm{W}_{O}^{s}$, we need first to ensure $\hat{\bm{W}}_{O}^{0}=\alpha\bm{W}_{O}^{0}$. Suppose that we have $\hat{\bm{W}}_{O}^{s}=\bm{W}_{O}^{s}$, then the update rule is:
$\hat{\bm{W}}_{O}^{s+1}$ $=\hat{\bm{W}}_{O}^{s}-\eta^{\prime}\nabla_{\hat{\bm{W}}_{O}^{s}}\mathcal{L}(\hat{\bm{W}}_{O}^{s})$ (34)
$=\alpha\bm{W}_{O}^{s}-\eta^{\prime}\nabla_{\bm{W}}\mathcal{L}(\bm{W})|_{\bm{W}=\alpha\bm{W}_{O}^{s}}\textrm{,}$ (35)
To ensure $\hat{\bm{W}}_{O}^{s+1}=\alpha\bm{W}_{O}^{s+1}$, we need the new learning rate $\eta^{\prime}$ meets $\eta^{\prime}=\alpha^{2}\eta$. For advanced optimization algorithms, e.g., Adam (Kingma & Ba, 2014) and AdamW (Loshchilov & Hutter, 2017). We have the following update rule (take AdamW for example, $\gamma$ refers to the weight decay ratio):
$\bm{g}_{s+1}$ $=\nabla_{\bm{W}_{O}^{s}}\mathcal{L}(\alpha\bm{W}_{O}^{s})=\alpha\nabla_{\bm{W}}\mathcal{L}(\bm{W})|_{\bm{W}=\alpha\bm{W}_{O}^{s}}$ (36)
$\bm{m}_{s+1}$ $=\beta_{1}\bm{m}_{s}+(1-\beta_{1})\bm{g}_{s+1}$ (37)
$\bm{v}_{s+1}$ $=\beta_{2}\bm{v}_{s}+(1-\beta_{2})\bm{g}^{2}_{s+1}$ (38)
$\bm{W}_{O}^{s+1}$ $=(1-\eta\gamma)\bm{W}_{O}^{s}-\eta\frac{\bm{m}_{s+1}/(1-\beta_{1}^{t})}{\sqrt{\bm{v}_{s+1}/(1-\beta_{2}^{t})}+\epsilon}$ (39)
We denote $\hat{\bm{g}}_{s+1},\,\hat{\bm{m}}_{s+1},\,\hat{\bm{v}}_{s+1}$ represents the intermediate counterparts for update of scenario where we only modify learning rate and initialization. First, we also need to ensure the initialization $\hat{\bm{W}}_{O}^{0}=\alpha\bm{W}_{O}^{0}$. Then we assume that we have already matched $\hat{\bm{W}}_{O}^{s}=\alpha\bm{W}_{O}^{s}$. The gradient for each step is
$\hat{\bm{g}}_{s+1}=\nabla_{\hat{\bm{W}}_{O}^{s}}\mathcal{L}(\hat{\bm{W}}_{O}^{s})=\nabla_{\bm{W}}\mathcal{L}(\bm{W})|_{\bm{W}=\alpha\bm{W}_{O}^{s}}=\bm{g}_{s+1}/\alpha$ (40)
Then the first and second moment will be $\hat{\bm{m}}_{s+1}=\bm{m}_{s+1}/\alpha$ and $\hat{\bm{v}}_{s+1}=\bm{v}_{s+1}/\alpha^{2}$. The updated weights will be
$\hat{\bm{W}}_{O}^{s+1}$ $=(1-\eta^{\prime}\gamma^{\prime})\hat{\bm{W}}_{O}^{s}-\eta^{\prime}\frac{\hat{\bm{m}}_{s+1}/(1-\beta_{1}^{t})}{\sqrt{\hat{\bm{v}}_{s+1}/(1-\beta_{2}^{t})}+\epsilon^{\prime}}$ (41)
$=(1-\eta^{\prime}\gamma^{\prime})\alpha\bm{W}_{O}^{s}-\eta^{\prime}\frac{\bm{m}_{s+1}/\alpha(1-\beta_{1}^{t})}{\sqrt{\bm{v}_{s+1}/\alpha^{2}(1-\beta_{2}^{t})}+\epsilon^{\prime}}$ (42)
Therefore, to ensure $\hat{\bm{W}}_{O}^{s+1}=\alpha\bm{W}_{O}^{s+1}$, one solution is $\eta^{\prime}=\alpha\eta$ and $\epsilon^{\prime}=\epsilon/\alpha$ and $\gamma^{\prime}=\gamma/\alpha$.
@@
Normalizer. Besides the normalizer $\bm{Z}_{i}=\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j^{\prime}}))$ considered in the main paper, we consider alternative normalizer: $\bm{Z}_{i}=\left(\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j^{\prime}}))^{p}\right)^{\frac{1}{p}}$, which gives us following attention operation:
$\bm{v}_{i}^{\dagger}=\frac{\sum_{j=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))\bm{v}_{j}}{\bm{Z}_{i}}=\frac{\sum_{j=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))\bm{v}_{j}}{\left(\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j^{\prime}}))^{p}\right)^{\frac{1}{p}}}\textrm{.}$ (43)
For softmax attention, we have $\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))=\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})$, then we can derive that
$\bm{v}_{i}^{\dagger}=\frac{\sum_{j=1}^{i}\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}}))\bm{v}_{j}}{\left(\sum_{j^{\prime}=1}^{i}\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}})^{p}\right)^{\frac{1}{p}}}=\sum_{j=1}^{i}\left(\frac{\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}/p})}{\sum_{j^{\prime}=1}^{i}\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}/p})}\right)^{\frac{1}{p}}\bm{v}_{j}\textrm{.}$ (44)
This is equivalent to adding a temperature $1/p$ into the softmax attention logits, and then taking the $p$-root of attention scores after softmax. We call this $p$-normalized softmax attention. $p=\textrm{1}$ in regular softmax attention. Similarly, we construct the LMs with $p$-normalized sigmoid attention.
We find that when $p=2$ or $p=3$ or $p=4$, pre-training of $p$-normalized softmax attention diverges and the loss goes infinity. When $p=1/2$ or $p=1/3$ or $p=1/4$, LM pre-training converges. As the attention scores are not added up to one, we investigate the massive activations instead, as visualized in Figure 29(Left). With smaller $p$, massive activations are mitigated to some extent, but not as effective as sigmoid attention without normalization. Intuitively, smaller $p$ induces a larger temperature in softmax operation, which leads to flattened attention logits.
Afterward, we conduct experiments on $p$-normalized sigmoid attention. There is no training problem with $p$ larger than 1. As visualized in Figure 29(Right), LMs with $p$-normalized sigmoid attention still demonstrate strong massive activations. To conclude, different normalizers may affect the amplitude of attention sink, but not stop its emergence.
Figure 29: $\ell_{2}$-norm ratio of $\bm{h}_{1}^{l}$ and mean of $\bm{h}_{t\neq 0}^{l}$. (Left) $p$-normalized softmax attention and also sigmoid attention without normalization (as reference). (Right) $p$-normalized sigmoid attention and also sigmoid attention without normalization (as reference).
@@
$\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}})$ 18.18 3.73
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})$ $1$       0.44 3.70
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}})$ 30.24 3.74
$\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})+1$ $1$       0.80 3.69
$\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})+1$ $\sum_{j^{\prime}=1}^{i}\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}})+1$
$\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j})+1)^{\top}}{\sqrt{d_{h}}}$ $\sum_{j^{\prime}=1}^{i}\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j^{\prime}})+1)^{\top}}{\sqrt{d_{h}}}$    53.65 4.19
$\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j})+1)^{\top}}{\sqrt{d_{h}}}$ $1$
$\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}}$ $\textrm{max}\left(\left|\sum_{j^{\prime}=1}^{i}\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}}\right|\textrm{,}\,1\right)$
$\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}}$ $1$    0.00 3.99
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{\top}}{\sqrt{d_{h}}}$ $\textrm{max}\left(\left|\sum_{j^{\prime}=1}^{i}\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j^{\prime}})^{\top}}{\sqrt{d_{h}}}\right|\textrm{,}\,1\right)$    0.19 3.85
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{\top}}{\sqrt{d_{h}}}$ $1$    0.74 3.91
Appendix E More experiments in LM after pre-training
Training stability in supervised fine-tuning. To investigate the long-term impacts of attention sink on model behaviors after pre-training, we conduct supervised fine-tuning (SFT) on our pre-trained 1B LMs with softmax attention and sigmoid attention without normalization. Specifically, we utilize the platform [^1] https://github.com/huggingface/alignment-handbook to conduct our experiments. The experimental configurations include: the UltraChat dataset (about 200k training samples) [^2] https://huggingface.co/datasets/HuggingFaceH4/ultrachat_200k (Ding et al., 2023), full-model fine-tuning, a learning rate of 2e-5 with cosine scheduling, batch size of 64, each of which contains 2048 tokens, one training epoch. As shown in Figure 30, we monitor the training loss and gradient norm of our two LMs during SFT. These two models behave similarly in terms of the above two metrics. Additionally, despite no attention sink, LMs with sigmoid attention without normalization have no issues of training stability during SFT. Though not from the attention sink perspective, a concurrent work by Ramapuram et al. (2024) discussed the theory and practices for Transformer models with sigmoid attention in detail. We refer the readers to Ramapuram et al. (2024) for more analyses.
Figure 30: The training loss and gradient norm in our 1B LMs with softmax attention and sigmoid attention without normalization in supervised fine-tuning.